# Unlasting: Unpaired Single-Cell Multi-Perturbation Estimation by Dual Conditional Diffusion Implicit Bridges

### v1

@@ 1 Introduction
Different single-cell perturbations, including CRISPR-based gene knockouts [2, 14] and small-molecule treatments [20], act at different layers of cellular mechanisms. Despite significant advancements in sequencing technology, producing perturbation data remains costly and time-consuming. As it is impractical to perform experiments across all cell types and perturbation conditions, accurately predicting perturbation responses under novel conditions is crucial. This capability significantly enhances biomedical research, particularly in advancing the understanding of gene functions and accelerating drug screening.
RNA-seq requires cell lysis to release RNA during sequencing, making it an irreversible and destructive process for cells [18]. Consequently, in single-cell perturbation experiments, capturing the same cell’s phenotype before and after perturbation is not feasible (Fig. 1). As a result, single-cell perturbation data are fundamentally unpaired. Although existing methods [21, 13, 3, 28, 11, 7] for predicting cell responses under unseen perturbation conditions have made significant progress, they often overlook the inherently unpaired nature of single-cell perturbation data, either by forcibly matching samples from the perturbed and unperturbed groups or by disregarding their relationships during modeling. On the other hand, while the unpaired nature of the data has been considered in some studies [5, 6], their use of unconditional models prevents them from generalizing to novel perturbation settings.
To address these issues, we propose Unlasting (Unpaired Single-Cell Multi-Perturbation Estimation by Dual Conditional Diffusion Implicit Bridges), a method leverages Dual Diffusion Implicit Bridges (DDIB,[25]) to predict single-cell responses to unseen genetic and molecular perturbations. Unlasting primarily consists of two parts: the source model and the target model. The source model learns the distribution of the unperturbed group, while the target model learns the distribution of the perturbed group. Both models share the same prior space, allowing it to establish a bridge between the unperturbed and perturbed states without requiring explicit pairing of samples. Besides, our model incorporates gene regulatory network (GRN) information to provide biologically meaningful guidance during perturbation modeling, improving the interpretability of cellular responses to perturbations. Given the sparsity of gene expression, we design a mask model to predict silent genes, thereby improving the quality of the generated profiles. Moreover, we observe that some genes exhibit bimodal expression under the same condition, indicating substantial heterogeneity in single-cell responses. To better capture this, we propose a more suitable evaluation metric beyond expectation-based assessments.
@@
Figure 2: Overview of Unlasting. Unlasting leverages DDIB [25] to predict cellular responses under unseen perturbation conditions. The source model obtain the latent embedding $x^{l}$ by adding DDIM-based forward noise to unperturbed cell sample $x^{c}$. Then, conditioned on the perturbation, we apply DDIM denoising to $x^{t}$ to generate the predicted sample.
@@ 2.1 Gene Regulation Network Construction and Molecule Representation Extraction
Gene regulatory networks (GRNs) describe gene interactions within a cell, but existing networks rely on manual annotations and are limited by cell types, hindering generalization. To address this, foundation models [8, 10, 30] have emerged to automatically learn universal gene regulatory patterns from large datasets. Our dataset enables more efficient extraction of reliable GRN structures. Additionally, advances in unsupervised molecular representation methods [31, 29] allow the extraction of features from unlabeled chemical data, capturing patterns in small molecules. This progress allows for more accurate modeling of the effects of small molecule drugs on cells.
@@ 2.2 Perturbation Estimation Model
Genetic and molecular perturbations constitute the two main research directions in single-cell perturbation studies. Existing methods have made significant progress in modeling single-cell perturbation responses. Some approaches rely on graph-based regression models to predict the outcomes of perturbations [21, 7]. Other methods employ generative models to reconstruct the distribution of perturbed states [16, 8, 12, 28, 3]. However, many of these approaches largely overlook the intrinsic relationship between control and perturbed samples during modeling. A separate class of methods enforces explicit pairing between unperturbed and perturbed samples, which may introduce unrealistic assumptions about the data.
@@ 2.3 Diffusion Process
In this section, we introduce the basic formulation of diffusion [17, 9]. Given an input sample $x_{0}$, we progressively add noise to it via the forward diffusion process as follows:
@@ 2.4 DDIM Inversion
The DDIM ([22]) proposes a straightforward inversion technique based on the ODE process, which significantly accelerates the inversion of $x_{T}$ back to $x_{0}$, based on the assumption that the ODE process can be reversed in the limit of small steps, , which can be written as:
@@ 3 Methodology
In this section, we introduce the proposed model Unlasting. The overview is shown in Fig. 2. Specifically, the source model learns the distribution of unperturbed cells, while the target model learns the distribution of cells under various perturbation conditions. By using a source model and a target model that share a prior space, we align the distributions of unperturbed and perturbed cells, thereby addressing the issue of unpaired data. Furthermore, in Section. 3.6, we provide a new interpretation of the effectiveness of the DDIB [25], viewing it as a form of data augmentation.
@@ 3.2 Data Preprocessing and Gene Regulation Network Construction
We first apply the SCANPY package [27] to perform log1p normalization on the gene expression data, and then select the top $N$ highly variable genes (HVGs). To facilitate stable training, we normalize the gene expression values to the range $[0,1]$ using the max value $x_{max}$ from the test set after splitting the dataset as: $x^{\prime}=\frac{x}{x_{\max}}$. When generating predictions, we restore the normalized values back to the original scale by multiplying by $x_{max}$.
When initializing the gene regulatory network (GRN), we first use the pre-trained foundation model [8] to obtain a basic GRN $\bar{A}\in R^{N\times N}$. However, the vocabulary of the foundation model may not include all of our target genes. Therefore, we supplement the $\bar{A}$ using co-expression information. Specifically, for a pair of genes $i$ and $j$, if the absolute value of their Pearson correlation coefficient (PCC) exceeds a given threshold $\epsilon_{co}$, we set $A_{i,j}=1$.
@@ 3.3 Conditional Diffusion Model
Unlike traditional diffusion models [17], which predict noise at a certain time step (Eq. 2), gene expression data presents a unique challenge due to the complex and less structured nature of the noise, making its modeling significantly more difficult. Therefore, our model directly predicts $x_{0}$, the clean gene expression data. The model outputs can be uniformly written as:
@@ 3.4 Gene Regulation Network based Block
Similarly, when $\mathbb{P}=\tt{mole}$, the perturbation condition $P=\{\mathbb{S},\mathbb{D}\}$, where $\mathbb{S}\in R^{D_{\mathbb{S}}}$ denotes the representation of the drug molecule extracted by the pre-trained molecular model [31], and $\mathbb{D}\in R$ represents the drug dose. These representations are then fused together through an MLP, $\Psi$, to obtain a combined perturbation condition embedding $F_{\mathbb{S,D}}=\Psi(\mathbb{S},\mathbb{D})\in R^{D}$.
@@
In the case of $\mathbb{P}=\tt{gene}$, the perturbation condition is given by $P={k}$, which biologically corresponds to the knockout of a specific gene $k$. We incorporate this perturbation information as follow:
@@
where $H$ represents the number of head, and $F^{0}$ is initialized as the embedding matrix $G_{\mathbb{P}}$, as defined earlier. The $\operatorname{GAT}$ used for feature aggregation can be found in [4, 26]. The final output of the GAT layers is $\widetilde{G}_{\mathbb{P}}=[\widetilde{g}_{\mathbb{P}}^{1},\widetilde{g}_{\mathbb{P}}^{2},\dots,\widetilde{g}_{\mathbb{P}}^{N}]$. Finally, we obtain the embedding $F_{GW}\in R^{N}$, which contains gene-wise information (Eq. 12). This embedding summarizes the perturbation effects for each individual gene and is passed to other model modules for downstream processing.
@@ 3.5 Implementation and Generation
In predicting the perturbation results, we adopt DDIM [22], which uses an ODE-based process. We first add noise to the unperturbed cell gene expression sample $x^{c}$, obtaining its latent embedding $x^{l}$ (Eq. 14. a). During denoising, we use the real sample $x^{c}$ from the unperturbed group in place of $\mu_{c},\sigma_{c}$ in Eq. 5, and generate the prediction $x^{t}$ under perturbation condition $P$ (Eq. 14. b).
[…]
@@ 4 Experiments and Results
In the main experiments, we use the Adamson [1] dataset of CRISPR knockouts and sci-Plex3 [24] dataset of chemical perturbations. Adamson contains data from 87 types of single-gene perturbations, with a single cell type. sci-Plex3 consists of 187 perturbation drugs, with four different dosage levels, and the cell come from three distinct cell types. In both datasets, each condition combination is observed in an average of over 100 cells. We consider $5,000$ genes in the Adamson dataset and $2,000$ genes in sci-Plex3 dataset.
@@ 4.1 Experiment Settings and Bimodal Expression Characteristics
In the training process, we randomly select 70% of gene perturbation conditions for the training set and use the remaining for testing in the Adamson dataset. In the SciPlex3 dataset, we first designate all samples under certain drug conditions [23, 12] as the OOD (Out-of-Distribution) test set. For the remaining samples, we randomly select samples from certain dosage levels under each drug-cell type condition as the test set, while the rest are used for training. The number of head in Eq. 11 is set to $2$. The $\epsilon_{co}$ in Eq. 11 is $0.3$ in both datasets. The batch size for model training is set to $32$, and the diffusion process is configured with a total of $500$ steps. For inference, we adopt DDIM sampling with $50$ steps to accelerate generation while maintaining sample quality. For datasets Adamson and SciPlex3, training steps are adjusted to $20,000$ and $100,000$, respectively. All our method and its competitors are conducted using one Nvidia A100 GPU.
@@ 4.2 Unlasting outperform existing methods
In this section, We compare our model with several baseline methods to evaluate its effectiveness in predicting gene expression under perturbations. These include: CPA[15], GEARS [21], GraphVCI [28], scGPT [8], chemCPA [12] and GRAPE [7].
Table 2 shows that Unlasting outperforms [21, 8, 7, 28], which rely on forced pairing of perturbed and unperturbed cells during training. This reliance on paired data limits their ability to capture true cellular heterogeneity, causing these models to converge towards average effects and miss the full diversity of cellular responses. Moreover, the suboptimal performance of [28] is also attributed to its insufficient modeling of the semantic meaning of perturbation conditions. In contrast, Unlasting explicitly incorporates a GRN block to more faithfully model the biological effects of perturbations. Methods like [12] and [15] further underperform because they reconstruct only perturbed cells without modeling the transition from the unperturbed state, and they assume gene expression follows a Gaussian distribution, which poorly reflects reality (see Fig. 5). Crucially, Unlasting overcomes the limitations of paired data by employing dual implicit bridges to explicitly and flexibly model the relationship between unperturbed and perturbed states, enabling more accurate and biologically faithful predictions.
@@ 4.3 Unlasting Performs Well on OOD Drug Perturbation and Double Gene Perturbation
To further validate the effectiveness of Unlasting, we evaluate its performance on double gene knockouts using the Norman dataset ([19]) and on out-of-distribution (OOD) drugs, as described in Section 4.1. Double gene knockouts involve complex gene–gene interactions, and experimental results show that our model effectively captures these interactions. To predict the effects of double gene perturbations, we use all observed samples under single gene perturbations and unperturbed conditions as the training set. OOD drugs, which are not seen during training, primarily target epigenetic regulation, tyrosine kinase signaling, and cell cycle regulation [23]. These drugs are representative of key biological processes and are often distinct from the drug in the training set. Our model demonstrates superior performance, suggesting that it better captures the effects of unseen molecules on cellular behavior.

### v2

@@ 1 Introduction
Different single-cell perturbations, including CRISPR-based gene knockouts Barrangou and Doudna (2016); Lino et al. (2018) and small-molecule treatments Peidli et al. (2024), act at different layers of cellular mechanisms. Despite significant advancements in sequencing technology, producing perturbation data remains costly and time-consuming. As it is impractical to perform experiments across all cell types and perturbation conditions, accurately predicting perturbation responses under novel conditions is crucial. This capability significantly enhances biomedical research, particularly in advancing the understanding of gene functions and accelerating drug screening.
RNA-seq requires cell lysis to release RNA during sequencing, making it an irreversible and destructive process for cells Mortazavi et al. (2008). Consequently, in single-cell perturbation experiments, capturing the same cell’s phenotype before and after perturbation is not feasible (Fig. 1). As a result, single-cell perturbation data are fundamentally unpaired. Although existing methods Roohani et al. (2022); Hetzel et al. (2022b); Bereket and Karaletsos (2024); Wu et al. (2022); He et al. (2024); Chi et al. (2025) for predicting cell responses under unseen perturbation conditions have made significant progress, they often overlook the inherently unpaired nature of single-cell perturbation data, either by forcibly matching samples from the perturbed and unperturbed groups or by disregarding their relationships during modeling. On the other hand, while the unpaired nature of the data has been considered in some studies Bunne et al. (2023); Cao et al. (2024), their use of unconditional models prevents them from generalizing to novel perturbation settings.
To address these issues, we propose Unlasting (Unpaired Single-Cell Multi-Perturbation Estimation by Dual Conditional Diffusion Implicit Bridges), a method leverages Dual Diffusion Implicit Bridges (DDIB,Su et al. (2022)) to predict single-cell responses to unseen genetic and molecular perturbations. Unlasting primarily consists of two parts: the source model and the target model. The source model learns the distribution of the unperturbed group, while the target model learns the distribution of the perturbed group. Both models share the same prior space, allowing it to establish a bridge between the unperturbed and perturbed states without requiring explicit pairing of samples. Besides, our model incorporates gene regulatory network (GRN) information to provide biologically meaningful guidance during perturbation modeling, improving the interpretability of cellular responses to perturbations. Given the sparsity of gene expression, we design a mask model to predict silent genes, thereby improving the quality of the generated profiles. Moreover, we observe that some genes exhibit bimodal expression under the same condition, indicating substantial heterogeneity in single-cell responses. To better capture this, we propose a more suitable evaluation metric beyond expectation-based assessments.
@@
Figure 2: Overview of Unlasting. Unlasting leverages DDIB Su et al. (2022) to predict cellular responses under unseen perturbation conditions. The source model obtain the latent embedding $x^{l}$ by adding DDIM-based forward noise to unperturbed cell sample $x^{c}$. Then, conditioned on the perturbation, we apply DDIM denoising to $x^{t}$ to generate the predicted sample.
@@ 2.1 Gene Regulation Network Construction and Molecule Representation Extraction
Gene regulatory networks (GRNs) describe gene interactions within a cell, but existing networks rely on manual annotations and are limited by cell types, hindering generalization. To address this, foundation models Cui et al. (2024); Hao et al. (2024); Yang et al. (2024) have emerged to automatically learn universal gene regulatory patterns from large datasets. Our dataset enables more efficient extraction of reliable GRN structures. Additionally, advances in unsupervised molecular representation methods Zhou et al. (2023); Xia et al. (2023) allow the extraction of features from unlabeled chemical data, capturing patterns in small molecules. This progress allows for more accurate modeling of the effects of small molecule drugs on cells.
@@ 2.2 Perturbation Estimation Model
Genetic and molecular perturbations constitute the two main research directions in single-cell perturbation studies. Existing methods have made significant progress in modeling single-cell perturbation responses. Some approaches rely on graph-based regression models to predict the outcomes of perturbations Roohani et al. (2022); Chi et al. (2025). Other methods employ generative models to reconstruct the distribution of perturbed states Lotfollahi et al. (2019); Cui et al. (2024); Hetzel et al. (2022a); Wu et al. (2022); Bereket and Karaletsos (2024). However, many of these approaches largely overlook the intrinsic relationship between control and perturbed samples during modeling. A separate class of methods enforces explicit pairing between unperturbed and perturbed samples, which may introduce unrealistic assumptions about the data.
@@ 2.3 Diffusion Process
In this section, we introduce the basic formulation of diffusion Luo (2022); Guo et al. (2023). Given an input sample $x_{0}$, we progressively add noise to it via the forward diffusion process as follows:
@@ 2.4 DDIM Inversion
The DDIM (Song et al. (2020)) proposes a straightforward inversion technique based on the ODE process, which significantly accelerates the inversion of $x_{T}$ back to $x_{0}$, based on the assumption that the ODE process can be reversed in the limit of small steps, , which can be written as:
@@ 3 Methodology
In this section, we introduce the proposed model Unlasting. The overview is shown in Fig. 2. Specifically, the source model learns the distribution of unperturbed cells, while the target model learns the distribution of cells under various perturbation conditions. By using a source model and a target model that share a prior space, we align the distributions of unperturbed and perturbed cells, thereby addressing the issue of unpaired data. Furthermore, in Section. 3.6, we provide a new interpretation of the effectiveness of the DDIB Su et al. (2022), viewing it as a form of data augmentation.
@@ 3.2 Data Preprocessing and Gene Regulation Network Construction
We first apply the SCANPY package Wolf et al. (2018) to perform log1p normalization on the gene expression data, and then select the top $N$ highly variable genes (HVGs). To facilitate stable training, we normalize the gene expression values to the range $[0,1]$ using the max value $x_{max}$ from the test set after splitting the dataset as: $x^{\prime}=\frac{x}{x_{\max}}$. When generating predictions, we restore the normalized values back to the original scale by multiplying by $x_{max}$.
When initializing the gene regulatory network (GRN), we first use the pre-trained foundation model Cui et al. (2024) to obtain a basic GRN $\bar{A}\in R^{N\times N}$. However, the vocabulary of the foundation model may not include all of our target genes. Therefore, we supplement the $\bar{A}$ using co-expression information. Specifically, for a pair of genes $i$ and $j$, if the absolute value of their Pearson correlation coefficient (PCC) exceeds a given threshold $\epsilon_{co}$, we set $A_{i,j}=1$.
@@ 3.3 Conditional Diffusion Model
Unlike traditional diffusion models Luo (2022), which predict noise at a certain time step (Eq. 2), gene expression data presents a unique challenge due to the complex and less structured nature of the noise, making its modeling significantly more difficult. Therefore, our model directly predicts $x_{0}$, the clean gene expression data. The model outputs can be uniformly written as:
@@ 3.4 Gene Regulation Network based Block
Similarly, when $\mathbb{P}=\tt{mole}$, the perturbation condition $P=\{\mathbb{S},\mathbb{D}\}$, where $\mathbb{S}\in R^{D_{\mathbb{S}}}$ denotes the representation of the drug molecule extracted by the pre-trained molecular model Zhou et al. (2023), and $\mathbb{D}\in R$ represents the drug dose. These representations are then fused together through an MLP, $\Psi$, to obtain a combined perturbation condition embedding $F_{\mathbb{S,D}}=\Psi(\mathbb{S},\mathbb{D})\in R^{D}$.
@@
In the case of $\mathbb{P}=\tt{gene}$, the perturbation condition is given by P=kP={k}, which biologically corresponds to the knockout of a specific gene $k$. We incorporate this perturbation information as follow:
@@
where $H$ represents the number of head, and $F^{0}$ is initialized as the embedding matrix $G_{\mathbb{P}}$, as defined earlier. The $\operatorname{GAT}$ used for feature aggregation can be found in Brody et al. (2021); Veličković et al. (2017). The final output of the GAT layers is $\widetilde{G}_{\mathbb{P}}=[\widetilde{g}_{\mathbb{P}}^{1},\widetilde{g}_{\mathbb{P}}^{2},\dots,\widetilde{g}_{\mathbb{P}}^{N}]$. Finally, we obtain the embedding $F_{GW}\in R^{N}$, which contains gene-wise information (Eq. 12). This embedding summarizes the perturbation effects for each individual gene and is passed to other model modules for downstream processing.
@@ 3.5 Implementation and Generation
In predicting the perturbation results, we adopt DDIM Song et al. (2020), which uses an ODE-based process. We first add noise to the unperturbed cell gene expression sample $x^{c}$, obtaining its latent embedding $x^{l}$ (Eq. 14. a). During denoising, we use the real sample $x^{c}$ from the unperturbed group in place of $\mu_{c},\sigma_{c}$ in Eq. 5, and generate the prediction $x^{t}$ under perturbation condition $P$ (Eq. 14. b).
$x^{l}$ $=\tt{ODESolve}(\hat{x}_{\theta},x^{c},c,0,1)$ $\quad\text{(a)}\quad x^{t}$ $=\tt{ODESolve}(\hat{x}_{\theta},x^{l},c,x^{c},P,1,0)$ (b) (14)
@@ 4 Experiments and Results
In the main experiments, we use the Adamson Adamson et al. (2016) dataset of CRISPR knockouts and sci-Plex3 Srivatsan et al. (2020b) dataset of chemical perturbations. Adamson contains data from 87 types of single-gene perturbations, with a single cell type. sci-Plex3 consists of 187 perturbation drugs, with four different dosage levels, and the cell come from three distinct cell types. In both datasets, each condition combination is observed in an average of over 100 cells. We consider $5,000$ genes in the Adamson dataset and $2,000$ genes in sci-Plex3 dataset.
@@ 4.1 Experiment Settings and Bimodal Expression Characteristics
In the training process, we randomly select 70% of gene perturbation conditions for the training set and use the remaining for testing in the Adamson dataset. In the SciPlex3 dataset, we first designate all samples under certain drug conditions Srivatsan et al. (2020a); Hetzel et al. (2022a) as the OOD (Out-of-Distribution) test set. For the remaining samples, we randomly select samples from certain dosage levels under each drug-cell type condition as the test set, while the rest are used for training. The number of head in Eq. 11 is set to $2$. The $\epsilon_{co}$ in Eq. 11 is $0.3$ in both datasets. The batch size for model training is set to $32$, and the diffusion process is configured with a total of $500$ steps. For inference, we adopt DDIM sampling with $50$ steps to accelerate generation while maintaining sample quality. For datasets Adamson and SciPlex3, training steps are adjusted to $20,000$ and $100,000$, respectively. All our method and its competitors are conducted using one Nvidia A100 GPU.
@@ 4.2 Unlasting outperform existing methods
In this section, We compare our model with several baseline methods to evaluate its effectiveness in predicting gene expression under perturbations. These include: CPALotfollahi et al. , GEARS Roohani et al. (2022), GraphVCI Wu et al. (2022), scGPT Cui et al. (2024), chemCPA Hetzel et al. (2022a) and GRAPE Chi et al. (2025).
Table 2 shows that Unlasting outperforms Roohani et al. (2022); Cui et al. (2024); Chi et al. (2025); Wu et al. (2022), which rely on forced pairing of perturbed and unperturbed cells during training. This reliance on paired data limits their ability to capture true cellular heterogeneity, causing these models to converge towards average effects and miss the full diversity of cellular responses. Moreover, the suboptimal performance of Wu et al. (2022) is also attributed to its insufficient modeling of the semantic meaning of perturbation conditions. In contrast, Unlasting explicitly incorporates a GRN block to more faithfully model the biological effects of perturbations. Methods like Hetzel et al. (2022a) and Lotfollahi et al. further underperform because they reconstruct only perturbed cells without modeling the transition from the unperturbed state, and they assume gene expression follows a Gaussian distribution, which poorly reflects reality (see Fig. 5). Crucially, Unlasting overcomes the limitations of paired data by employing dual implicit bridges to explicitly and flexibly model the relationship between unperturbed and perturbed states, enabling more accurate and biologically faithful predictions.
@@ 4.3 Unlasting Performs Well on OOD Drug Perturbation and Double Gene Perturbation
To further validate the effectiveness of Unlasting, we evaluate its performance on double gene knockouts using the Norman dataset (Norman et al. (2019)) and on out-of-distribution (OOD) drugs, as described in Section 4.1. Double gene knockouts involve complex gene–gene interactions, and experimental results show that our model effectively captures these interactions. To predict the effects of double gene perturbations, we use all observed samples under single gene perturbations and unperturbed conditions as the training set. OOD drugs, which are not seen during training, primarily target epigenetic regulation, tyrosine kinase signaling, and cell cycle regulation Srivatsan et al. (2020a). These drugs are representative of key biological processes and are often distinct from the drug in the training set. Our model demonstrates superior performance, suggesting that it better captures the effects of unseen molecules on cellular behavior.